OpenStudy (anonymous):

would an equation of a line with a slope of 4 and a y-intercept of (o,2) be y=4x+2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

one last question for you, youve been such a big help

OpenStudy (anonymous):

whats the equation of a line with a slope of 3/7 and contains the point (3,8)

OpenStudy (anonymous):

okay back to the line equation y = mx + b given slope so plug it in y = (3/7)x + b now plug in the points 8 = (3/7)3 + b b = 8 - 9/7 = 47/7 so y = (3/7)x + 47/7

OpenStudy (anonymous):

find the equation of a line that contains the points (3,7) and (12, -4)

OpenStudy (anonymous):

lols you said last question try this one, i'll walk you through it what do you have to do first?

OpenStudy (anonymous):

y=mx+b form?

OpenStudy (anonymous):

wait, find the slope first

OpenStudy (anonymous):

which would be -11/9

OpenStudy (anonymous):

yup to both how do you find the slope of two points

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

now you have the formula y = mx + b and you have the slope and two points, now what?

OpenStudy (anonymous):

so as of right now the equation is y=-11/9x+b. you put in the two points in the equation?

OpenStudy (anonymous):

since both points are on the line you can choose either ONE pick one to put into the equation it doens't matter which one you use, you'll get the same answer

OpenStudy (anonymous):

so i could have 7=(-11/9)(3)+b?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and then you solve for b?

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

can you walk me through solving b one more time?

OpenStudy (anonymous):

sure you have 7=(-11/9)(3)+b multiply the fraction 7 = -11/3 + b add -11/3 on both sides b = 7 + 11/3 add b = 32/3

OpenStudy (anonymous):

when i multiplied the fraction i got -3.66. how did you get 11/3?

OpenStudy (anonymous):

nevermind. thats the fraction of -3.66

OpenStudy (anonymous):

lols you're using a calculator? -3.66 is the same as -11/3 put that into your calculator

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so final equation should be y=-11/3+32/3?

OpenStudy (anonymous):

you forgot an x y=(-11/3)x+32/3 ^-- here

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

cheers ^_^